# Layerwise Recurrent Router for Mixture-of-Experts

### v1

@@ Abstract
The scaling of large language models (LLMs) has revolutionized their capabilities in various tasks, yet this growth must be matched with efficient computational strategies. The Mixture-of-Experts (MoE) architecture stands out for its ability to scale model size without significantly increasing training costs. Despite their advantages, current MoE models often display parameter inefficiency. For instance, a pre-trained MoE-based LLM with 52 billion parameters might perform comparably to a standard model with 6.7 billion parameters (Rajbhandari et al., 2022). Being a crucial part of MoE, current routers in different layers independently assign tokens without leveraging historical routing information, potentially leading to suboptimal token-expert combinations and the parameter inefficiency problem. To alleviate this issue, we introduce the Layerwise Recurrent Router for Mixture-of-Experts (RMoE). RMoE leverages a Gated Recurrent Unit (GRU) to establish dependencies between routing decisions across consecutive layers. Such layerwise recurrence can be efficiently parallelly computed for input tokens and introduces negotiable costs. Our extensive empirical evaluations demonstrate that RMoE-based language models consistently outperform a spectrum of baseline models. Furthermore, RMoE integrates a novel computation stage orthogonal to existing methods, allowing seamless compatibility with other MoE architectures. Our analyses attribute RMoE’s gains to its effective cross-layer information sharing, which also improves expert selection and diversity. Our code is at https://github.com/qiuzh20/RMoE
@@ 1 Introduction
We validate RMoE’s performance with various model sizes, architectures, datasets, and training settings (per-training and supervised fine-tuning), demonstrating that RMoE outperforms a range of baselines. Moreover, RMoE’s introduction of a novel computation stage during routing makes it orthogonal to and compatible with most existing methods. We further analyze RMoE and elucidate the primary contributors to its improvement. Our findings indicate that while the relatively complex GRU in RMoE shares essential cross-layer information, it also enables additional gradient propagation for the router. Our analysis shows that layerwise recurrence provides cross-layer information, fostering router exploration and optimizing expert utilization. Consequently, the selected experts are leveraged more effectively, leading to increased diversity of experts. We believe that our innovative router design and massive analysis can offer insights into the development of future MoE models.
@@ 3.1 Preliminaries
The Gated Recurrent Units Dey & Salem (2017) module is an advanced variant of RNNs that addresses some of the traditional RNNs’ limitations, such as difficulty capturing long-term dependencies and gradient vanishing issues. Given an input $\mathbf{x}^{\prime}_{i}$ at time step $i$, GRU first calculates the reset gate $\mathbf{s}_{i}$ and the update gate $\mathbf{z}_{i}$ to determine how much of the previous memory to keep and to forget,
$\mathbf{s}_{i}$ $=\sigma(\mathbf{W}_{s}\mathbf{x}^{\prime}_{i}+\mathbf{U}_{s}\mathbf{h}_{i-1})$ (2)
$\mathbf{z}_{i}$ $=\sigma(\mathbf{W}_{z}\mathbf{x}^{\prime}_{i}+\mathbf{U}_{z}\mathbf{h}_{i-1})$
@@
$\mathbf{\tilde{h}}_{i}$ $=\tanh(\mathbf{W}_{h}\mathbf{x}^{\prime}_{i}+\mathbf{s}_{i}\odot(\mathbf{W}_{h}\mathbf{h}_{i-1}))$ (3)
$\mathbf{h}_{i}$ $=(1-\mathbf{z}_{i})\odot\mathbf{\tilde{h}}_{i}+\mathbf{z}_{i}\odot\mathbf{h}_{i-1}$
@@ 3.2 Layerwise Recurrent Router
Existing routers work independently, this lack of global information may prevent routers from discovering more effective token-expert combinations. Therefore, we integrate a GRU into the routing process, explicitly incorporating historical routing information into the current expert selection for each token. Formally, at the $i$ th layer, we first project the hidden state $\mathbf{x}_{i}$ to the dimension of the GRU state $\mathbf{x}^{\prime}_{i}\in\mathbb{R}^{p}$ (usually smaller than the dimension $h$ of $\mathbf{x}_{i}$. We choose 128 for most of the settings to save computational cost and provide further analysis in Tab. 6 and Tab. 7):
@@
Next, $\mathbf{h}_{i}$ is input into the router to make router decisions and then to aggregate different experts based on that decision:
@@ 4.1 Experimental settings
We mainly compare RMoE with other existing router designs. All methods are based on the decoder-only standard switch-transformer architecture. Following Pham et al. (2024), all routers adopt the same MoE configuration: selecting top-2 experts from 16 experts. Each task is trained on 2 NVIDIA A100 GPUs for about 20 hours. More training configurations can be found in App. A.2. Our baselines include (1) SMoE: standard switch-transformers with a standard linear router. (2) HyperMoE (Do et al., 2023): the method employs a fixed, randomly initialized hypernetwork (Ha et al., 2016) to produce the weights for the linear router, subsequently allowing the generated linear layer to perform the routing. (3) SMoE-MLP (Shen et al., 2023): it replaces the linear router with a two-layer MLP using the GELU activation function. (4) RandomMoE: inspired by SMoE-Dropout (Chen et al., 2023) and HyperMoE, we propose to compare with a fixed randomly initialized linear router; this could be a naive baseline for all learnable routers. (5) XMoE (Chi et al., 2022): it first down-projects the token embeddings to a lower dimension (default 16) and computes its cosine-similarity with the low-dimension expert embeddings. It also uses a learnable temperature in softmax operation. (5) CosineSMoE, similar to XMoE except without down-projection.
@@
As pre-training-then-supervised-fine-tuning has become the standard paradigm, we also evaluate the RMoE in this setting. Our pre-training corpus is a multilingual data collection that spans common and specialized domains, including Wikipedia, finance, and legal texts. We follow the filtering guidelines in Redpajama V2 (Computer, 2023) and subsequently perform a de-duplication process on the filtered data, yielding approximately 126.5 billion high-quality tokens. Our model architecture is modified based on Llama family (Touvron et al., 2023). Specifically, we use a 24-layer model and top-4 gating from 16 experts per layer following (Dai et al., 2024). This yields a model with approximately 0.53B activated / 0.91B total parameters. All different routers use the same training configurations. To ensure expert load balance, we employ balance loss with weights $0.01$ during training. These experiments are conducted using the Megablocks (Gale et al., 2023) on 8 NVIDIA A100 GPUs for about 5 days. More details can be found in App. A.2. After pertaining, we perform supervised fine-tuning (sft). All models are trained on Alpaca (Taori et al., 2023) with the same configuration. We use lm-evaluation-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness to evaluate the fine-tuned model. To simulate the real LLMs application scenario, we don’t perform task-specific fine-tuning and evaluation. Since the models are largely under-trained due to computation resource limitation, they give almost random-guessing results on challenging tasks like MMLU (Hendrycks et al., 2020). Therefore, we only test on tasks (ARC-easy, Hellaswag, PIQA, SciQ, LAMBADA) in lm-evaluation-harness, where the results largely surpass random guessing. More details about sft configurations and tasks can be found in App. A.2.
@@ 5 Ablation Studies
Following the aforementioned analysis, we try to further disentangle the effect of the layerwise recurrence. When removing the layerwise recurrence as in the RMoE + NP setting, we remove two information flows across layers: (1) the forward information about previous routers’ decisions and (2) the backward gradient propagation through GRU hidden states in different layers. To compare the two information flows, we investigate the following settings: (1) RMoE + detach $h_{i-1}$: in intermediate stage between RMoE and RMoE-NP. By detaching $h_{i-1}$ to stop its gradient computation in Eq. 5, each GRU cell can only use previous information during feed-forward. (2) RMoE + NP + r-$\alpha$: inspired by Realformer (He et al., 2020) that introduces residual attention score to facilitate attention gradient back-propagation, we investigate an intermediate stage between RMoE and RMoE-NP by adding gating logits residual for the RMoE + NP settings. Concretely, the gating score of $i$-th layer for expert $n$ is $g_{n}(\mathbf{h}_{i};\mathbf{G}_{i},k)+\alpha g_{n}(\mathbf{h}_{i-1};\mathbf{G}_{i-1},k)$. It is a straightforward way to supplement router information across layers based on the NP setting. In our experiments, we set $\alpha$ as 0.5 and 1.0. (3) Moreover, we also test detaching the gradient computation of passed logits ($\textbf{h}_{i-1}\times\textbf{G}_{i-1}$), denoted as ‘detach-r’. From Tab. 5, RMoE + detach $h_{i-1}$ performs even worse than RMoE-NP, showing that the Recurrent Gradient is important. Similarly, ‘NP+r0.5’ and ‘NP+r1.0’ are comparable with ‘NP’, demonstrating that the naive gating score residual cannot provide effective cross-layer information. The performance of their detached version largely drops, further demonstrating the importance of extra backward gradient passing.
Figure 2: Test BPC on Enwiki8 with different model sizes (6, 12, 18, 24, 32). Similar validation results are in App. A.3 Fig. 6
@@
This part tests the other components in RMoE, such as recurrent hidden state dimension, layerwise projector, and GRU cell. As shown in Tab. 6: (1) All methods with recurrent routers outperform SMoE, further validating the effectiveness of RMoE. (2) Layerwise projector in Eq. 5 performs better than standard RNNs using a single shared projector. One possible reason is that the weights and hidden states norm in different layers vary greatly (as shown in App. A.4.4 Fig. 13), and it would be hard for a single shared projector to process them. (3) The GRU router performs better than the RNN and LSTM routers. Moreover, we further compare RMoE variants in the larger scale settings. We compare pre-trained models with different structures and recurrent hidden dimensions in Tab. 7 (Averaged results, full results in App. A.3 Tab. 9). We can find similar results: (1) All RMoE variants outperform SMoE; (2) Simple router (RNN) and complex routers (GRU with $p=256,512$) perform worse than GRU with $p=128$. In short, layerwise projector and moderate recurrent cell (e.g. GRU with $p=128$) effectively introduce layerwise recurrent.
@@ 6 Observations
The intuition of the proposed RMoE is that current routers in different layers are isolated, and the layerwise GRU is incorporated to provide routers with some global information for coordination. Therefore, we measure the Mutual Information (MI) between routing distributions in different layers for each router. The code implementation can be found in App. A.4.1. We can observe: (1) Besides RMoE, all existing methods show low cross-layer MI, indicating that the routers of different layers work relatively independently. (2) RMoE shows higher MI than three baselines (d v.s. a, b, and c) and RMoE-NP (d v.s. e), showing the recurrent router can facilitate cross-layer information sharing. (3) While RMoE-NP’s MI is largely smaller than RMoE, it still surpasses the three baseline methods. The reason can be the shared GRU in Eq. 5. (4) Intuitively, passing routing logits can directly improve MI (f v.s. e). However, directly passing logits can’t ensure long-range information sharing, as the values in the right part of (f), which indicate the MI between non-neighbor layers, are smaller than those in (d).
@@
Figure 4: Gate score entropy distribution over Enwiki8 test set for different router configurations. More similar results can be found in App. A.4.2 Fig. 9 and Fig. 10.
@@
We also look into the statistics of selected experts’ scores. Here we calculate the (1) Inner Balance (IB): defined as the ratio $\operatorname{Top-1}$ score / $\operatorname{Top-k}$ score, large IB means the first expert dominates all selected experts; and (2) Outer Balance (OB), defined as $\sum_{k\in\operatorname{Top-k}}g_{k}$, indicating the selected scores’ ratio in the score distribution, large OB means selected expert scores dominate the gate score distribution. Because such a ratio could have some extreme values, we report the median number for all tokens in Tab. 8. We can observe: (1) RandomMoE, with a fixed router, shows the lowest IB and OB. (2) Low-entropy models in the previous section (Sec. 6) have high IB and OB. (3) RMoE gives suitable IB and OB. While simply using a complex router (‘RMoE-NP’) shows relatively low IB and OB, RMoE is even lower. Moreover, passing logits can reduce IB and OB (‘RMoE-NP-r0.5/1.0’). All these experiments show sharing cross-layer router information can lead to more balanced routing decision and thus facilitate expert usage.
@@ 7 Conclusion
8 Limitations
Our pre-trained language model is still under-trained and not tested on popular challenging tasks like MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021). Introducing moderate z-loss (Zoph et al., 2022) (penalizes large router logits) and changing temperature in softmax operation (leads to more flat score distributions) to achieve moderate flat gating scores. We will investigate these modifications and combine them with RMoE in future works. The supplement gradient pathways hypothesis is only investigated by indirect observations like performance (Sec 5). More direct gradient analysis for MoE’s router might bring more insights.

### v2

@@ Abstract
The scaling of large language models (LLMs) has revolutionized their capabilities in various tasks, yet this growth must be matched with efficient computational strategies. The Mixture-of-Experts (MoE) architecture stands out for its ability to scale model size without significantly increasing training costs. Despite their advantages, current MoE models often display parameter inefficiency. For instance, a pre-trained MoE-based LLM with 52 billion parameters might perform comparably to a standard model with 6.7 billion parameters (Rajbhandari et al., 2022). Being a crucial part of MoE, current routers in different layers independently assign tokens without leveraging historical routing information, potentially leading to suboptimal token-expert combinations and the parameter inefficiency problem. To alleviate this issue, we introduce the Layerwise Recurrent Router for Mixture-of-Experts (RMoE). RMoE leverages a Gated Recurrent Unit (GRU) to establish dependencies between routing decisions across consecutive layers. Such layerwise recurrence can be efficiently parallelly computed for input tokens and introduces negotiable costs. Our extensive empirical evaluations demonstrate that RMoE-based language models consistently outperform a spectrum of baseline models. Furthermore, RMoE integrates a novel computation stage orthogonal to existing methods, allowing seamless compatibility with other MoE architectures. Our analyses attribute RMoE’s gains to its effective cross-layer information sharing, which also improves expert selection and diversity. Our code is at https://github.com/qiuzh20/RMoE.
@@ 1 Introduction
We validate RMoE’s performance with various model sizes, architectures, datasets, and training settings (per-training and supervised fine-tuning), demonstrating that RMoE outperforms a range of baselines. Moreover, RMoE’s introduction of a novel computation stage during routing makes it orthogonal to and compatible with most existing methods. We further analyze RMoE and elucidate the primary contributors to its improvement. Our findings indicate that while the GRU in RMoE shares essential cross-layer information, it also enables additional gradient propagation for the router. Our analysis shows that layerwise recurrence provides cross-layer information, fostering router exploration and optimizing expert utilization. Consequently, the selected experts are leveraged more effectively, leading to increased diversity of experts. We believe that our innovative router design and massive analysis can offer insights into the development of future MoE models.
@@ 3.1 Preliminaries
The Gated Recurrent Units (GRU) Dey & Salem (2017) module is an advanced variant of RNNs that addresses traditional RNNs’ limitations, such as difficulty capturing long-term dependencies and gradient vanishing issues. Given an input $\mathbf{x}^{\prime}_{i}$ at time step $i$, GRU first calculates the reset gate $\mathbf{s}_{i}$ and the update gate $\mathbf{z}_{i}$ to determine how much of the previous memory to keep and to forget,
$\mathbf{s}_{i}$ $=\sigma(\mathbf{W}_{s}\mathbf{x}^{\prime}_{i}+\mathbf{U}_{s}\mathbf{h}_{i-1}),\,\,\,\,\,\,\,\,\mathbf{z}_{i}=\sigma(\mathbf{W}_{z}\mathbf{x}^{\prime}_{i}+\mathbf{U}_{z}\mathbf{h}_{i-1})$ (2)
@@
$\mathbf{\tilde{h}}_{i}$ $=\tanh(\mathbf{W}_{h}\mathbf{x}^{\prime}_{i}+\mathbf{s}_{i}\odot(\mathbf{W}_{h}\mathbf{h}_{i-1})),\,\,\,\,\,\,\,\,\mathbf{h}_{i}=(1-\mathbf{z}_{i})\odot\mathbf{\tilde{h}}_{i}+\mathbf{z}_{i}\odot\mathbf{h}_{i-1}$ (3)
@@ 3.2 Layerwise Recurrent Router
Existing routers work independently, this lack of global information may prevent routers from discovering more effective token-expert combinations. Therefore, we integrate a GRU into the routing process, explicitly incorporating historical routing information into the current expert selection for each token. Formally, at the $i$ th layer, we first use a linear layer to project the hidden state $\mathbf{x}_{i}$ to the dimension of the GRU state $\mathbf{x}^{\prime}_{i}\in\mathbb{R}^{p}$ (usually smaller than the dimension $h$ of $\mathbf{x}_{i}$. We choose 128 for most of the settings provide further analysis in Tab. 6 and Tab. 7):
@@
Next, $\mathbf{h}_{i}$ is input into the router and then expert outputs are aggregated based on the router output:
@@ 4.1 Experimental settings
We compare RMoE with other existing router designs. All methods are based on the decoder-only standard switch-transformer architecture with post-norm. Following (Pham et al., 2024), all routers select top-2 experts from 16 experts. Each task is trained on 2 NVIDIA A100 GPUs for about 20 hours. More training configurations can be found in App. A.2. Our baselines include (1) SMoE: standard switch-transformers with a standard linear router. (2) HyperMoE (Do et al., 2023): the method employs a fixed, randomly initialized hypernetwork (Ha et al., 2016) to produce the weights for the linear router, subsequently allowing the generated linear layer to perform the routing. (3) SMoE-MLP (Shen et al., 2023): it replaces the linear router with a two-layer MLP using the GELU activation function. (4) RandomMoE: inspired by SMoE-Dropout (Chen et al., 2023) and HyperMoE, we propose to compare with a fixed randomly initialized linear router; this could be a naive baseline for all learnable routers. (5) XMoE (Chi et al., 2022): it first down-projects the token embeddings to a lower dimension (default 16) and computes its cosine-similarity with the low-dimension expert embeddings. It also uses a learnable temperature in softmax. (5) CosineSMoE, similar to XMoE except without down-projection.
@@
As pre-training-then-supervised-fine-tuning has become the standard paradigm, we also evaluate the RMoE in this setting. We conduct preliminary scale-up experiment on a setting of training 0.91B models with 40B tokens. Our pre-training corpus is a multilingual data collection that spans common and specialized domains, including Wikipedia, finance, and legal texts. Our model architecture is modified based on Llama family (Touvron et al., 2023). Specifically, we use a 24-layer model and top-4 gating from 16 experts per layer following (Dai et al., 2024). This yields a model with approximately 0.53B activated / 0.91B total parameters. All different routers use the same training configurations. To ensure expert load balance, we employ balance loss with weights $0.01$ during training. These experiments are conducted using the Megablocks (Gale et al., 2023) on 8 NVIDIA A100 GPUs for about 5 days. More details can be found in App. A.2. After pertaining, we perform supervised fine-tuning (sft). All models are trained on Alpaca (Taori et al., 2023) with the same configuration. We use lm-evaluation-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness to evaluate the fine-tuned model. To simulate the real LLMs application scenario, we don’t perform task-specific fine-tuning and evaluation. Since the models are largely under-trained, they give almost random-guessing results on challenging tasks like MMLU (Hendrycks et al., 2020). Therefore, we only test on tasks (ARC-easy, Hellaswag, PIQA, SciQ, LAMBADA) in lm-evaluation-harness. More details about sft configurations and tasks can be found in App. A.2. We further justify the scalability of RMoE on the setting of training 15B activate 2.7B models with 120B / 400B tokens. Given our utilization of a high-quality pre-training corpus, pre-training on 400B tokens yields better results compared to experimental MoE like OpenMoE (Xue et al., 2024). We find RMoE consistently provides over a one-point improvement in performance on benchmarks such as MMLU, GSM8K, and HumanEval. More details can be found in App. A.3.
@@ 5 Ablation Studies
Following the aforementioned analysis, we try to further disentangle the effect of the layerwise recurrence. When removing the layerwise recurrence as in the RMoE + NP setting, we remove two information flows across layers: (1) the forward information about previous routers’ decisions and (2) the backward gradient propagation through GRU hidden states in different layers. To compare the two information flows, we investigate the following settings: (1) RMoE + detach $h_{i-1}$: in intermediate stage between RMoE and RMoE-NP. By detaching $h_{i-1}$ to stop its gradient computation in Eq. 5, each GRU cell can only use previous information during feed-forward. (2) RMoE + NP + r-$\alpha$: inspired by Realformer (He et al., 2020) that introduces residual attention score to facilitate attention gradient back-propagation, we investigate an intermediate stage between RMoE and RMoE-NP by adding gating logits residual for the RMoE + NP settings. Concretely, the gating score of $i$-th layer for expert $n$ is $g_{n}(\mathbf{h}_{i};\mathbf{G}_{i},k)+\alpha g_{n}(\mathbf{h}_{i-1};\mathbf{G}_{i-1},k)$. It is a straightforward way to supplement router information across layers based on the NP setting. In our experiments, we set $\alpha$ as 0.5 and 1.0. (3) Moreover, we also test detaching the gradient computation of passed logits ($\textbf{h}_{i-1}\times\textbf{G}_{i-1}$), denoted as ‘detach-r’. From Tab. 5, RMoE + detach $h_{i-1}$ performs even worse than RMoE-NP, showing that the Recurrent Gradient is important. Similarly, ‘NP+r0.5’ and ‘NP+r1.0’ are comparable with ‘NP’, showing that the naive gating score residual can’t provide effective cross-layer information. The performance of their detached version largely drops, demonstrating the importance of extra gradient passing.
Figure 2: Test BPC on Enwiki8 with different model sizes (6, 12, 18, 24, 32). Similar validation results are in App. A.5 Fig. 14
@@
This part tests the other components in RMoE, such as recurrent hidden state dimension, layerwise projector, and GRU cell. As shown in Tab. 6: (1) All methods with recurrent routers outperform SMoE. (2) Layerwise projector in Eq. 5 performs better than standard RNNs using a single shared projector. One possible reason is that the weights and hidden states norm in different layers vary greatly (as shown in App. A.4.5 Fig. 11), and it would be hard for a single shared projector to process them. This approach aligns with the design principle of not sharing LayerNorm parameters when employing shared MoE transformer blocks, as discussed by Xue et al. (2022). (3) The GRU router performs best. Moreover, we further compare RMoE variants in the larger scale settings. We compare pre-trained models with different structures and recurrent hidden dimensions in Tab. 7 (Averaged results, full results in App. A.5 Tab. 12). We can find similar results: (1) All RMoE variants outperform SMoE; (2) Simple router (RNN) and complex routers (GRU with $p=256,512$) perform worse. In short, layerwise projector and moderate recurrent cell (e.g. GRU with $p=128$) effectively introduce layerwise recurrent.
@@ 6 Observations
The intuition of the proposed RMoE is that current routers in different layers are isolated, and the layerwise GRU is incorporated to provide routers with global information for coordination. Therefore, we measure the Mutual Information (MI) between routing distributions in different layers for each router in Fig. 3. The code can be found in App. A.4.2. We can observe: (1) Besides RMoE, all existing methods show low cross-layer MI, indicating that the routers of different layers work relatively independently. (2) RMoE shows higher MI than three baselines (d v.s. a, b, and c) and RMoE-NP (d v.s. e), showing the recurrent router can facilitate cross-layer information sharing. (3) While RMoE-NP’s MI is largely smaller than RMoE, it still surpasses the three baseline methods. The reason can be the shared GRU in Eq. 5. (4) Intuitively, passing routing logits can directly improve MI (f v.s. e). However, directly passing logits can’t ensure long-range information sharing, as the values in the right part of (f), which indicate the MI between non-neighbor layers, are smaller than those in (d).
@@
Figure 4: Gate score entropy distribution over Enwiki8 test set for different router configurations. More similar results can be found in App. A.4.4 Fig. 8 and Fig. 9.
@@
We also look into the statistics of selected experts’ scores. Here we calculate the (1) Inner Balance (IB): defined as the ratio $\operatorname{Top-1}$ score/$\operatorname{Top-2}$ score, large IB means the first expert dominates all selected experts; and (2) Outer Balance (OB), defined as $\sum_{k\in\operatorname{Top-k}}g_{k}$, indicating the selected scores’ ratio in the score distribution, large OB means selected expert scores dominate the gate score distribution. Because such a ratio could have some extreme values, we report the median number for all tokens in Tab. 8. We can observe: (1) RandomMoE, with a fixed router, shows the lowest IB and OB. (2) Low-entropy models in the previous section (Sec. 6) have high IB and OB. (3) RMoE gives suitable IB and OB. While simply using a complex router (‘RMoE-NP’) shows relatively low IB and OB, RMoE is even lower. Moreover, passing logits can reduce IB and OB (‘RMoE+NP+r-0.5/1.0’). All these experiments show sharing cross-layer router information can lead to more balanced routing decision and thus facilitate expert usage.
Layerwise recurrence reduces the negative effect of load balance constraint.
To provide a more direct analysis of the router gradient, we investigate how the gradient norm of the router varies throughout the entire training process. When training a MoE model, the gradient of the router has two separate sources: (1) the language modeling (LM) loss, and (2) the load balancing (LB) loss that pushes the router to assign tokens to different experts in a balanced manner. We empirically find (1) LB loss dominates the training of the linear router at the early training stage. This could hurt model’s general performance, as Wang et al. (2024) find, a high LB loss can cause balance token distribution but reduce performance. (2) On the contrary, the gradient of the RNN router from LB loss stabilises in the early stage, and the gradient from the LM loss keeps decreasing, suggesting that the RNN router is more optimised towards the LM loss. These observations suggest the recurrent router can effectively controls the influence of the LB loss. More details can be found in App. A.4.1
@@ 7 Conclusion
Acknowledgment
Jie Fu is supported by Shanghai Artificial Intelligence Laboratory.